OpenStudy (thivitaa):

x³-3x²-2=0 factorize please?

OpenStudy (jim766):

x^3 - 3x^2 -2 = 0 (x^3 + 1) + (-3x^2 + 3) do you see what I have done here? just pulled them apart creating a sum of 2 cubes in the 1st

OpenStudy (jim766):

(x^3 + 1) + (-3x^2 + 3) (x+ 1)(x^2 -x + 1) +(-3x^2 + 3) sum of 2 cubes

OpenStudy (jim766):

(x^3 + 1) + (-3x^2 + 3) (x+ 1)(x^2 -x + 1) +(-3x^2 + 3) sum of 2 cubes (x+ 1)(x^2 -x + 1) -3(x^2 - 1) factor out the -3 you with me?

OpenStudy (thivitaa):

owh okk @Jim766 thanks !!!

OpenStudy (jim766):

(x^3 + 1) + (-3x^2 + 3) (x+ 1)(x^2 -x + 1) +(-3x^2 + 3) sum of 2 cubes (x+ 1)(x^2 -x + 1) -3(x^2 - 1) factor out the -3 (x+1)(x^2-x+1) -3(x+ 1)(x -1)= 0 difference of 2 squares

OpenStudy (jim766):

(x+1)(x^2-x+1) -3(x+ 1)(x -1)= 0 difference of 2 squares (x+1) [(x^2 -x+ 1) -3(x-1)] factor out the (x+1) (x+1)(x^2-x+1 -3x + 3) you still with me?

OpenStudy (thivitaa):

ya @Jim766 getting it in my head... thanks tho..:)

OpenStudy (jim766):

(x+1)(x^2-x+1 -3x + 3) = 0 (x+1)(x^2 -4x + 4) = 0 (x + 2)(x -2)^2 = 0 (x + 1)( x - 2)(x - 2) = 0 x + 1= 0 , x -2 = 0 solutions are -1 and 2